• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name A. Camsonne

    Log entry time 20:37:31 on February 3,2009

    Entry number 258814

    keyword=Final tune for Compton, rates are good

    Rates are back to about 150 Hz/uA I will hand back the beam to the shift crew to go back to production.
    Turning back laser on

    FIGURE 1